# The numeraire e-variable

### 29 Feb 2024 v1

@@ 2 The numeraire e-variable
Consider the space $L^{0}_{+}({\mathbb{Q}})$ consisting of all nonnegative real-valued random variables modulo ${\mathbb{Q}}$-almost sure equivalence. Let $C\subset L^{0}_{+}({\mathbb{Q}})$ be the set of (equivalence classes of) random variables that are ${\mathbb{Q}}$-almost surely equal to a finite e-variable. The set $C$ is convex, solid, and closed in probability. Suppose now that ${\mathcal{E}}$ is not bounded in probability under ${\mathbb{Q}}$. Then $C$ is not bounded in probability. The decomposition result in Lemma 2.3 of Brannath and Schachermayer (1999) then shows that there is an event $A$ with ${\mathbb{Q}}(A)>0$ such that $n\bm{1}_{A}$ belongs to $C$ for all $n\in{\mathbb{N}}$. By definition of $C$, for each $n$ there is a finite e-variable $X_{n}$ that is ${\mathbb{Q}}$-almost surely equal to $n\bm{1}_{A}$, say on an event $B_{n}$ with ${\mathbb{Q}}(B_{n})=1$. By Fatou’s lemma, $X=\liminf_{n\to\infty}X_{n}$ is again an e-variable. Moreover, $X=\infty$ on $A\cap(\bigcap_{n\in\mathbb{N}}B_{n})$, which coincides with $A$ up to a ${\mathbb{Q}}$-nullset and thus has positive probability under ${\mathbb{Q}}$. In other words, $X$ is an e-variable that is not ${\mathbb{Q}}$-almost surely finite. ∎
The following result, usually referred to as a Komlós-type lemma, is the basis of the proof that a numeraire exists. Here $\operatorname{conv}(\ldots)$ denotes the convex hull of a set of random variables, that is, the collection of finite convex combinations of its elements. For a proof, see Lemma A1.1 and the subsequent Remark 1 in Delbaen and Schachermayer (1994).
@@
Let $(X_{n})_{n\in{\mathbb{N}}}$ be any sequence of $[0,\infty]$-valued random variables. One can choose $\widetilde{X}_{n}\in\operatorname{conv}(X_{n},X_{n+1},\ldots)$ such that $\widetilde{X}_{n}$ converges ${\mathbb{Q}}$-almost surely to a $[0,\infty]$-valued random variable.
@@
Let $U$ be a utility function that is bounded from above. There exists an e-variable $X^{*}$ that attains $\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{\mathbb{Q}}[U(X)]$ and has the property that every e-variable is ${\mathbb{Q}}$-almost surely finite on $\{X^{*}<\infty\}$. If ${\mathbb{E}}_{\mathbb{Q}}[U^{\prime}(X^{*})X^{*}]<\infty$, one has the first-order condition
${\mathbb{E}}_{\mathbb{Q}}\left[U^{\prime}(X^{*})(X-X^{*})\right]\leq 0,\quad X\in{\mathcal{E}}.$ (2.2)
@@
Let $X_{n}\in{\mathcal{E}}$ form a sequence such that ${\mathbb{E}}_{\mathbb{Q}}[U(X_{n})]$ increases to the supremum. Lemma 2.7 yields a ${\mathbb{Q}}$-almost surely convergent sequence of random variables of the form $\widetilde{X}_{n}=\sum_{i=n}^{m_{n}}\lambda^{n}_{i}X_{i}$ for convex weights $\lambda^{n}_{n},\ldots,\lambda^{n}_{m_{n}}$. Define $X^{*}=\liminf_{n\to\infty}\widetilde{X}_{n}$ and note that while this may be a proper limit inferior for some $\omega\in\Omega$, we have ${\mathbb{Q}}$-almost surely that $X^{*}=\lim_{n\to\infty}\widetilde{X}_{n}$. Each $\widetilde{X}_{n}$ belongs to ${\mathcal{E}}$ since this set is convex, and Fatou’s lemma then implies that $X^{*}$ belongs to ${\mathcal{E}}$ as well. Next, since $U$ is concave and ${\mathbb{E}}[U(X_{n})]$ is an increasing sequence,
${\mathbb{E}}_{\mathbb{Q}}[U(\widetilde{X}_{n})]\geq\sum_{i=n}^{m_{n}}\lambda^{n}_{i}{\mathbb{E}}_{\mathbb{Q}}[U(X_{i})]\geq{\mathbb{E}}_{\mathbb{Q}}[U(X_{n})].$ (2.3)
[…]
${\mathbb{E}}_{\mathbb{Q}}[U(X^{*})]\geq\limsup_{n\to\infty}{\mathbb{E}}_{\mathbb{Q}}[U(\widetilde{X}_{n})]\geq\limsup_{n\to\infty}{\mathbb{E}}_{\mathbb{Q}}[U(X_{n})]=\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{\mathbb{Q}}[U(X)].$
@@
We next show that $X^{*}$ can be chosen so that every e-variable is ${\mathbb{Q}}$-almost surely finite on $\{X^{*}<\infty\}$. This uses Lemma A.1, which in particular states that there is an event $N$ that is ${\mathcal{P}}$-negligible, meaning that ${\mathbb{P}}(N)=0$ for all ${\mathbb{P}}\in{\mathcal{P}}$, and satisfies ${\mathbb{Q}}(N^{\prime}\setminus N)=0$ for any other ${\mathcal{P}}$-negligible event $N^{\prime}$. Define $X^{**}=X^{*}\bm{1}_{N^{c}}+\infty\bm{1}_{N}$. This is still an e-variable since $N$ is ${\mathcal{P}}$-negligible, and it attains the supremum since it is at least as large as $X^{*}$. Furthermore, since $N^{\prime}=\{X=\infty\}$ is ${\mathcal{P}}$-negligible for any $X\in{\mathcal{E}}$, we have (${\mathbb{Q}}(\{X=\infty\}\setminus\{X^{**}=\infty\})=0$. Equivalently, by taking complements, $X$ is ${\mathbb{Q}}$-almost surely finite on $\{X^{**}<\infty\}$. Replacing $X^{*}$ by $X^{**}$ thus ensures the claimed property.
@@
${\mathbb{E}}_{\mathbb{Q}}\left[\frac{U(X(t))-U(X(0))}{t}\right]\leq 0,\quad t\in(0,1].$
@@
${\mathbb{E}}_{\mathbb{Q}}\left[U^{\prime}(X^{*})(\delta+(1-\delta)X-X^{*})\right]\leq 0.$
@@
and that $U_{n}^{\prime}$ is convex. Applying Lemma 2.8 with $U=U_{n}$ gives an optimal e-variable $X_{n}^{*}$ for each $n$. These have the additional property that the sets $\{X_{n}^{*}<\infty\}$ all coincide up to ${\mathbb{Q}}$-nullsets, say with a fixed set $A$. Because $U^{\prime}_{n}(x)x\leq 1$ for all $x\in(0,\infty]$, we may use the first-order condition (2.2) to deduce
${\mathbb{E}}_{\mathbb{Q}}\left[U_{n}^{\prime}(X_{n}^{*})X\bm{1}_{A}\right]\leq{\mathbb{Q}}(A),\quad X\in{\mathcal{E}}.$ (2.5)
Next, Lemma 2.7 yields a ${\mathbb{Q}}$-almost surely convergent sequence of random variables $\widetilde{X}_{n}^{*}=\sum_{i=n}^{m_{n}}\lambda^{n}_{i}X_{i}^{*}$ for some convex weights $\lambda^{n}_{n},\ldots,\lambda^{n}_{m_{n}}$. We define $X^{*}=\liminf_{n\to\infty}\widetilde{X}_{n}^{*}$, which belongs to ${\mathcal{E}}$ and is ${\mathbb{Q}}$-almost surely the limit of $\widetilde{X}_{n}^{*}$. Note that $A$ then coincides ${\mathbb{Q}}$-almost surely with the sets where $X^{*}$ and the $\widetilde{X}^{*}_{n}$ are finite. Using first (2.5), then (2.4), and finally the convexity of $U^{\prime}_{n}$ we get
${\mathbb{Q}}(A)\geq{\mathbb{E}}_{\mathbb{Q}}\left[\sum_{i=n}^{m_{n}}\lambda^{n}_{i}U_{i}^{\prime}(X_{i}^{*})X\bm{1}_{A}\right]\geq{\mathbb{E}}_{\mathbb{Q}}\left[\sum_{i=n}^{m_{n}}\lambda^{n}_{i}U_{n}^{\prime}(X_{i}^{*})X\bm{1}_{A}\right]\geq{\mathbb{E}}_{\mathbb{Q}}\left[U_{n}^{\prime}(\widetilde{X}_{n}^{*})X\bm{1}_{A}\right].$
The expression inside the last expectation is nonnegative and converges ${\mathbb{Q}}$-almost surely to $X\bm{1}_{A}/X^{*}$. In particular, $X^{*}$ is ${\mathbb{Q}}$-almost surely strictly positive on $A$. Fatou’s lemma now yields ${\mathbb{E}}_{\mathbb{Q}}[X\bm{1}_{A}/X^{*}]\leq{\mathbb{Q}}(A)$. On $A^{c}$, where $X^{*}=\infty$, we have $X/X^{*}\leq 1$ thanks to the convention (1.1). Hence ${\mathbb{E}}_{\mathbb{Q}}[X/X^{*}]\leq{\mathbb{E}}_{\mathbb{Q}}[X\bm{1}_{A}/X^{*}]+{\mathbb{Q}}(A^{c})\leq 1$. This shows that $X^{*}$ is a numeraire. ∎
@@ 3 Duality and reverse information projection
We continue to consider an arbitrary nonempty composite null hypothesis ${\mathcal{P}}$ and a simple alternative ${\mathbb{Q}}$ on a measurable space $(\Omega,{\mathcal{F}})$. There is a duality relationship between ${\mathcal{P}}$ and its set of e-variables. To develop this perspective, recall that $F_{+}$ and $M_{+}$ denote the set of all $[0,\infty]$-valued measurable functions and nonnegative measures on $\Omega$, respectively. The polar of ${\mathcal{P}}$ under the pairing $(X,{\mathbb{P}})\mapsto{\mathbb{E}}_{\mathbb{P}}[X]$ is defined by
${\mathcal{P}}^{\circ}=\{X\in F_{+}\colon{\mathbb{E}}_{\mathbb{P}}[X]\leq 1\text{ for all }{\mathbb{P}}\in{\mathcal{P}}\}.$
This is precisely the set ${\mathcal{E}}$ of e-variables, i.e., ${\mathcal{P}}^{\circ}={\mathcal{E}}$. The bipolar of ${\mathcal{P}}$ is
${\mathcal{P}}^{\circ\circ}=\{{\mathbb{P}}\in M_{+}\colon{\mathbb{E}}_{\mathbb{P}}[X]\leq 1\text{ for all }X\in{\mathcal{P}}^{\circ}\}.$
The bipolar can be interpreted as the effective null hypothesis: it is the largest family of distributions whose set of e-variables is exactly ${\mathcal{E}}$. Thus it consists of those distributions against which e-variables in ${\mathcal{E}}$ do not provide any evidence. Here ‘distribution’ should be understood as sub-probability measure. Because the constant $X=1$ is always an e-variable, elements of ${\mathcal{P}}^{\circ\circ}$ have total mass at most one, and sometimes strictly less than one. For example, if ${\mathbb{P}}$ belongs to ${\mathcal{P}}^{\circ\circ}$, then so does any ${\mathbb{P}}^{\prime}$ in $M_{+}$ that is dominated by ${\mathbb{P}}$ in the sense that ${\mathbb{P}}^{\prime}(A)\leq{\mathbb{P}}(A)$ for all events $A$.
@@
The numeraire e-variable $X^{*}$ gives rise to an element *${\mathbb{P}}^{*}\in{\mathcal{P}}^{\circ\circ}$ defined by $d{\mathbb{P}}^{*}/d{\mathbb{Q}}=1/X^{*}$, understood as zero on $\{X^{*}=\infty\}$. Note ${\mathbb{P}}^{*}\ll{\mathbb{Q}}$ by definition.
@@
The measure ${\mathbb{P}}^{*}$ is well-defined since $X^{*}$ is ${\mathbb{Q}}$-almost surely positive. It belongs to ${\mathcal{P}}^{\circ\circ}$ because the numeraire property yields ${\mathbb{E}}_{{\mathbb{P}}^{*}}[X]={\mathbb{E}}_{\mathbb{Q}}[\bm{1}_{\{X^{*}<\infty\}}X/X^{*}]\leq{\mathbb{E}}_{\mathbb{Q}}[X/X^{*}]\leq 1$ for all $X\in{\mathcal{E}}={\mathcal{P}}^{\circ}$. ∎
Theorem 3.4 below along with results in Section 4 show that ${\mathbb{P}}^{*}$ can be interpreted as a reverse information projection, as we explain after that theorem. This concept was first introduced in the PhD thesis of Li (1999, Chapter 4), appeared centrally in the context of e-variables in Grünwald et al. (2024), and was developed further in a recent preprint by Lardy et al. (2023). This earlier work assumes the existence of a reference measure, an assumption we are able to dispense with but nonetheless examine further in Section 4.
@@
We refer to the measure ${\mathbb{P}}^{*}$ in Lemma 3.1 as the reverse information projection (RIPr) of ${\mathbb{Q}}$ on ${\mathcal{P}}^{\circ\circ}$.
@@
Our use of the RIPr terminology for ${\mathbb{P}}^{*}$ above is justified by what follows. Recall that the entropy of ${\mathbb{Q}}$ relative to ${\mathbb{P}}$ is defined by
@@
$H({\mathbb{Q}}\mid{\mathbb{P}})=\begin{cases}{\mathbb{E}}_{\mathbb{Q}}\left[\log\dfrac{d{\mathbb{Q}}}{d{\mathbb{P}}}\right]&\text{if }{\mathbb{Q}}\ll{\mathbb{P}},\\ +\infty&\text{otherwise.}\end{cases}$
@@
We will show that ${\mathbb{P}}^{*}$ minimizes the relative entropy $H({\mathbb{Q}}\mid{\mathbb{P}})$ over ${\mathcal{P}}^{\circ\circ}$. For this it will be convenient to use a slightly different expression for the relative entropy. For any ${\mathbb{P}}\in M_{+}$ we let ${\mathbb{P}}^{a}$ denote its absolutely continuous part with respect to ${\mathbb{Q}}$. One then has

### 22 Mar 2024 v2

@@ 2 The numeraire e-variable
Consider the space $L^{0}_{+}({\mathsf{Q}})$ consisting of all nonnegative real-valued random variables modulo ${\mathsf{Q}}$-almost sure equivalence. Let $C\subset L^{0}_{+}({\mathsf{Q}})$ be the set of (equivalence classes of) random variables that are ${\mathsf{Q}}$-almost surely equal to a finite e-variable. The set $C$ is convex, solid, and closed in probability. Suppose now that ${\mathcal{E}}$ is not bounded in probability under ${\mathsf{Q}}$. Then $C$ is not bounded in probability. The decomposition result in Lemma 2.3 of Brannath and Schachermayer (1999) then shows that there is an event $A$ with ${\mathsf{Q}}(A)>0$ such that $n\bm{1}_{A}$ belongs to $C$ for all $n\in{\mathbb{N}}$. By definition of $C$, for each $n$ there is a finite e-variable $X_{n}$ that is ${\mathsf{Q}}$-almost surely equal to $n\bm{1}_{A}$, say on an event $B_{n}$ with ${\mathsf{Q}}(B_{n})=1$. By Fatou’s lemma, $X=\liminf_{n\to\infty}X_{n}$ is again an e-variable. Moreover, $X=\infty$ on $A\cap(\bigcap_{n\in\mathbb{N}}B_{n})$, which coincides with $A$ up to a ${\mathsf{Q}}$-nullset and thus has positive probability under ${\mathsf{Q}}$. In other words, $X$ is an e-variable that is not ${\mathsf{Q}}$-almost surely finite. ∎
The following result, usually referred to as a ‘Komlós-type’ lemma, is the basis of the proof that a numeraire exists. For a proof, see Lemma A1.1 and the subsequent Remark 1 in Delbaen and Schachermayer (1994).
@@
Let $(X_{n})_{n\in{\mathbb{N}}}$ be any sequence of $[0,\infty]$-valued random variables. One can choose $\widetilde{X}_{n}\in\operatorname{conv}(X_{n},X_{n+1},\ldots)$ such that $\widetilde{X}_{n}$ converges ${\mathsf{Q}}$-almost surely to a $[0,\infty]$-valued random variable.
@@
Let $U$ be a utility function that is bounded from above. There exists an e-variable $X^{*}$ that attains $\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{\mathsf{Q}}[U(X)]$ and has the property that every e-variable is ${\mathsf{Q}}$-almost surely finite on $\{X^{*}<\infty\}$. If ${\mathbb{E}}_{\mathsf{Q}}[U^{\prime}(X^{*})X^{*}]<\infty$, one has the first-order condition
${\mathbb{E}}_{\mathsf{Q}}\left[U^{\prime}(X^{*})(X-X^{*})\right]\leq 0,\quad X\in{\mathcal{E}}.$ (2.2)
@@
Let $X_{n}\in{\mathcal{E}}$ form a sequence such that ${\mathbb{E}}_{\mathsf{Q}}[U(X_{n})]$ increases to the supremum. Lemma 2.7 yields a ${\mathsf{Q}}$-almost surely convergent sequence of random variables of the form $\widetilde{X}_{n}=\sum_{i=n}^{m_{n}}\lambda^{n}_{i}X_{i}$ for convex weights $\lambda^{n}_{n},\ldots,\lambda^{n}_{m_{n}}$. Define $X^{*}=\liminf_{n\to\infty}\widetilde{X}_{n}$ and note that while this may be a proper limit inferior for some $\omega\in\Omega$, we have ${\mathsf{Q}}$-almost surely that $X^{*}=\lim_{n\to\infty}\widetilde{X}_{n}$. Each $\widetilde{X}_{n}$ belongs to ${\mathcal{E}}$ since this set is convex, and Fatou’s lemma then implies that $X^{*}$ belongs to ${\mathcal{E}}$ as well. Next, since $U$ is concave and ${\mathbb{E}}[U(X_{n})]$ is an increasing sequence,
${\mathbb{E}}_{\mathsf{Q}}[U(\widetilde{X}_{n})]\geq\sum_{i=n}^{m_{n}}\lambda^{n}_{i}{\mathbb{E}}_{\mathsf{Q}}[U(X_{i})]\geq{\mathbb{E}}_{\mathsf{Q}}[U(X_{n})].$ (2.3)
Since $U$ is bounded above and $X^{*}=\lim_{n\to\infty}\widetilde{X}_{n}$ ${\mathsf{Q}}$-almost surely, we may use Fatou’s lemma together with (2.3) to get
${\mathbb{E}}_{\mathsf{Q}}[U(X^{*})]\geq\limsup_{n\to\infty}{\mathbb{E}}_{\mathsf{Q}}[U(\widetilde{X}_{n})]\geq\limsup_{n\to\infty}{\mathbb{E}}_{\mathsf{Q}}[U(X_{n})]=\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{\mathsf{Q}}[U(X)].$
@@
We next show that $X^{*}$ can be chosen so that every e-variable is ${\mathsf{Q}}$-almost surely finite on $\{X^{*}<\infty\}$. This uses Lemma A.1, which in particular states that there is an event $N$ that is ${\mathcal{P}}$-negligible, meaning that ${\mathsf{P}}(N)=0$ for all ${\mathsf{P}}\in{\mathcal{P}}$, and satisfies ${\mathsf{Q}}(N^{\prime}\setminus N)=0$ for any other ${\mathcal{P}}$-negligible event $N^{\prime}$. Define $X^{**}=X^{*}\bm{1}_{N^{c}}+\infty\bm{1}_{N}$. This is still an e-variable since $N$ is ${\mathcal{P}}$-negligible, and it attains the supremum since it is at least as large as $X^{*}$. Furthermore, since $N^{\prime}=\{X=\infty\}$ is ${\mathcal{P}}$-negligible for any $X\in{\mathcal{E}}$, we have ${\mathsf{Q}}(\{X=\infty\}\setminus\{X^{**}=\infty\})=0$. Equivalently, by taking complements, $X$ is ${\mathsf{Q}}$-almost surely finite on $\{X^{**}<\infty\}$. Replacing $X^{*}$ by $X^{**}$ thus ensures the claimed property.
@@
${\mathbb{E}}_{\mathsf{Q}}\left[\frac{U(X(t))-U(X(0))}{t}\right]\leq 0,\quad t\in(0,1].$
@@
${\mathbb{E}}_{\mathsf{Q}}\left[U^{\prime}(X^{*})(\delta+(1-\delta)X-X^{*})\right]\leq 0.$
@@
and that $U_{n}^{\prime}$ is convex. Applying Lemma 2.8 with $U=U_{n}$ gives an optimal e-variable $X_{n}^{*}$ for each $n$. These have the additional property that the sets $\{X_{n}^{*}<\infty\}$ all coincide up to ${\mathsf{Q}}$-nullsets, say with a fixed set $A$. Because $U^{\prime}_{n}(x)x\leq 1$ for all $x\in(0,\infty]$, we may use the first-order condition (2.2) to deduce
${\mathbb{E}}_{\mathsf{Q}}\left[U_{n}^{\prime}(X_{n}^{*})X\bm{1}_{A}\right]\leq{\mathsf{Q}}(A),\quad X\in{\mathcal{E}}.$ (2.5)
Next, Lemma 2.7 yields a ${\mathsf{Q}}$-almost surely convergent sequence of random variables $\widetilde{X}_{n}^{*}=\sum_{i=n}^{m_{n}}\lambda^{n}_{i}X_{i}^{*}$ for some convex weights $\lambda^{n}_{n},\ldots,\lambda^{n}_{m_{n}}$. We define $X^{*}=\liminf_{n\to\infty}\widetilde{X}_{n}^{*}$, which belongs to ${\mathcal{E}}$ and is ${\mathsf{Q}}$-almost surely the limit of $\widetilde{X}_{n}^{*}$. Note that $A$ then coincides ${\mathsf{Q}}$-almost surely with the sets where $X^{*}$ and the $\widetilde{X}^{*}_{n}$ are finite. Using first (2.5), then (2.4), and finally the convexity of $U^{\prime}_{n}$ we get
${\mathsf{Q}}(A)\geq{\mathbb{E}}_{\mathsf{Q}}\left[\sum_{i=n}^{m_{n}}\lambda^{n}_{i}U_{i}^{\prime}(X_{i}^{*})X\bm{1}_{A}\right]\geq{\mathbb{E}}_{\mathsf{Q}}\left[\sum_{i=n}^{m_{n}}\lambda^{n}_{i}U_{n}^{\prime}(X_{i}^{*})X\bm{1}_{A}\right]\geq{\mathbb{E}}_{\mathsf{Q}}\left[U_{n}^{\prime}(\widetilde{X}_{n}^{*})X\bm{1}_{A}\right].$
The expression inside the last expectation is nonnegative and converges ${\mathsf{Q}}$-almost surely to $X\bm{1}_{A}/X^{*}$. In particular, $X^{*}$ is ${\mathsf{Q}}$-almost surely strictly positive on $A$. Fatou’s lemma now yields ${\mathbb{E}}_{\mathsf{Q}}[X\bm{1}_{A}/X^{*}]\leq{\mathsf{Q}}(A)$. On $A^{c}$, where $X^{*}=\infty$, we have $X/X^{*}\leq 1$ thanks to the convention (1.1). Hence ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]\leq{\mathbb{E}}_{\mathsf{Q}}[X\bm{1}_{A}/X^{*}]+{\mathsf{Q}}(A^{c})\leq 1$. This shows that $X^{*}$ is a numeraire. ∎
@@ 3 Duality and reverse information projection
We continue to consider an arbitrary nonempty composite null hypothesis ${\mathcal{P}}$ and a simple alternative ${\mathsf{Q}}$ on a measurable space $(\Omega,{\mathcal{F}})$. There is a duality relationship between ${\mathcal{P}}$ and its set of e-variables. To develop this perspective, recall that $F_{+}$ and $M_{+}$ denote the set of all $[0,\infty]$-valued measurable functions and nonnegative measures on $\Omega$, respectively. The polar of ${\mathcal{P}}$ under the pairing $(X,{\mathsf{P}})\mapsto{\mathbb{E}}_{\mathsf{P}}[X]$ is defined by
${\mathcal{P}}^{\circ}=\{X\in F_{+}\colon{\mathbb{E}}_{\mathsf{P}}[X]\leq 1\text{ for all }{\mathsf{P}}\in{\mathcal{P}}\}.$
This is precisely the set of e-variables, i.e., ${\mathcal{P}}^{\circ}={\mathcal{E}}$. The bipolar of ${\mathcal{P}}$ is
${\mathcal{P}}^{\circ\circ}=\{{\mathsf{P}}\in M_{+}\colon{\mathbb{E}}_{\mathsf{P}}[X]\leq 1\text{ for all }X\in{\mathcal{P}}^{\circ}\}.$
The bipolar can be interpreted as the effective null hypothesis: it is the largest family of distributions whose set of e-variables is exactly ${\mathcal{E}}$. Thus it consists of those distributions against which e-variables in ${\mathcal{E}}$ do not provide any evidence. In particular, a ‘nontrivial’ e-variable $X$, in the sense that ${\mathbb{E}}_{\mathsf{Q}}[X]>1$, exists if and only if ${\mathsf{Q}}\notin{\mathcal{P}}^{\circ\circ}$. Because the constant $X=1$ is always an e-variable, elements of ${\mathcal{P}}^{\circ\circ}$ have total mass at most one, and sometimes strictly less than one. For example, if ${\mathsf{P}}$ belongs to ${\mathcal{P}}^{\circ\circ}$, then so does any ${\mathsf{P}}^{\prime}$ in $M_{+}$ that is dominated by ${\mathsf{P}}$ in the sense that ${\mathsf{P}}^{\prime}(A)\leq{\mathsf{P}}(A)$ for all $A$. In Section 4 we discuss more explicit descriptions of the bipolar; see Example 4.4, Lemma 4.5, and Corollary 4.6. A special case of the latter is Example 3.7 below, where the bipolar is computed for a simple null.
@@
The numeraire e-variable $X^{*}$ gives rise to an element ${\mathsf{P}}^{*}\in{\mathcal{P}}^{\circ\circ}$ defined by $d{\mathsf{P}}^{*}/d{\mathsf{Q}}=1/X^{*}$, understood as zero on $\{X^{*}=\infty\}$. Note that ${\mathsf{P}}^{*}\ll{\mathsf{Q}}$ by definition.
@@
Since $X^{*}$ is ${\mathsf{Q}}$-almost surely unique and strictly positive, ${\mathsf{P}}^{*}$ is well-defined. It belongs to ${\mathcal{P}}^{\circ\circ}$ because ${\mathbb{E}}_{{\mathsf{P}}^{*}}[X]={\mathbb{E}}_{\mathsf{Q}}[\bm{1}_{\{X^{*}<\infty\}}X/X^{*}]\leq{\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]\leq 1$ for all $X\in{\mathcal{E}}={\mathcal{P}}^{\circ}$ thanks to the numeraire property. ∎
Our results show that ${\mathsf{P}}^{*}$ can be interpreted as a reverse information projection, as we explain after Theorem 3.4 below. This concept was first introduced in the PhD thesis of Li (1999, Chapter 4), appeared centrally in the context of e-variables in Grünwald et al. (2024), and was developed further in a recent preprint by Lardy et al. (2023). This earlier work assumes the existence of a reference measure, an assumption we are able to dispense with but nonetheless examine further in Section 4.
@@
We refer to the measure ${\mathsf{P}}^{*}$ in Lemma 3.1 as the reverse information projection (RIPr) of ${\mathsf{Q}}$ on ${\mathcal{P}}^{\circ\circ}$.
@@
Our use of the RIPr terminology for ${\mathsf{P}}^{*}$ above is justified by what follows. Recall that the entropy of ${\mathsf{Q}}$ relative to ${\mathsf{P}}$ is defined by
@@
$H({\mathsf{Q}}\mid{\mathsf{P}})=\begin{cases}{\mathbb{E}}_{\mathsf{Q}}\left[\log\dfrac{d{\mathsf{Q}}}{d{\mathsf{P}}}\right]&\text{if }{\mathsf{Q}}\ll{\mathsf{P}},\\ +\infty&\text{otherwise.}\end{cases}$
@@
We will show that ${\mathsf{P}}^{*}$ minimizes the relative entropy $H({\mathsf{Q}}\mid{\mathsf{P}})$ over ${\mathcal{P}}^{\circ\circ}$. For this it will be convenient to use a slightly different expression for the relative entropy. For any ${\mathsf{P}}\in M_{+}$ we let ${\mathsf{P}}^{a}$ denote its absolutely continuous part with respect to ${\mathsf{Q}}$. One then has
$H({\mathsf{Q}}\mid{\mathsf{P}})={\mathbb{E}}_{\mathsf{Q}}\left[-\log\frac{d{\mathsf{P}}^{a}}{d{\mathsf{Q}}}\right].$ (3.1)
To see this, suppose first $d{\mathsf{P}}^{a}/d{\mathsf{Q}}$ is strictly positive ${\mathsf{Q}}$-almost surely. Then ${\mathsf{Q}}\ll{\mathsf{P}}^{a}\ll{\mathsf{P}}$ and $d{\mathsf{Q}}/d{\mathsf{P}}=d{\mathsf{Q}}/d{\mathsf{P}}^{a}=1/(d{\mathsf{P}}^{a}/d{\mathsf{Q}})$ ${\mathsf{Q}}$-almost surely, so (3.1) holds. If instead $d{\mathsf{P}}^{a}/d{\mathsf{Q}}$ is zero with positive ${\mathsf{Q}}$-probability, then the right-hand side of (3.1) is infinite (the expectation is always well-defined by Jensen’s inequality), and so is the relative entropy because ${\mathsf{Q}}\not\ll{\mathsf{P}}$. Thus (3.1) holds in this case, too.
Remark 3.3.
In the above, ${\mathsf{P}}$ can be any element of $M_{+}$. We do not adjust the definition of relative entropy for the possibility that its total mass is not one.
We now make the trivial observation that